if the payoff of rolling two dice is the sum of the upper face of the two dice, what is the expected payoff of one roll of the two dice?
Answer by Boreal(15235)

It is symmetric around the number 7.
(1/36)*2=2/36
(2/36)*3=6/36
(3/36)*4=12/36
(4/36)*5=20/36
(5/36)*6=30/36
(6/36)*7=42/36
(5/36)*8=40/36
(4/36)*9=36/36
(3/36)*10=30/36
(2/36)*11=22/36
(1/36)*12=12/36
sum of all of these values * probability of each is the expected value.
This adds to 252/36=7.